OpenStudy (anonymous):

Use Lagrange multipliers to find the dimensions of a box WITHOUT A TOP of volume V=1 of minimal surface area.

OpenStudy (anonymous):

Length * Height * Width= 1 Constraint -----> L H W =1 +G minimize this --> L W + 2 W H + 2 LH =F w.r.t <L, w, h> Gradient F <W+2H,L+2H, 2W+2L> Gradient of G <HW,LH,LW> \[\lambda*HW=W+2H\] \[\lambda*LH=L+2H\] \[\lambda*LW=2L+2W\]

OpenStudy (anonymous):

Thank you very, very much! Your work and response is extremely helpful!

OpenStudy (anonymous):

\[\lambda*HW=W+2H\] \[\lambda*LH=L+2H\] \[\lambda*LW=2L+2W\] \[L H W =1\] Four equations, four variable